OpenStudy (anonymous):

Can someone help me with Problem Set 1: 1B-13

OpenStudy (anonymous):

The question asks us to solve for cos(Θ1−Θ2), but the answer in the supplemental packet solves for cos(Θ2−Θ1). Is this a mistake?

OpenStudy (phi):

cos(-x) = cos(x) so their switching theta1 and theta2 does not change the answer. (Though it would be nice if they were consistent, so as not to introduce confusion)

OpenStudy (phi):

Here are more details on their derivation of the cos of a difference of two angles. First, they define a unit length vector with its tail at the origin and its head lying on the unit circle. Based on this definition, we know its i and j components  we thus have two unit vectors \[ \vec{u_1}= < \cos \theta_1 , \sin \theta_1> \\ \vec{u_2}= < \cos \theta_2 , \sin \theta_2> \]